A. What is the resistance of an electric frying pan that draws 12 amperes of current when connected to a 120-volt circuit?

 

B. How much current is drawn by a lamp that has a resistance of 100 ohms when a voltage of 50 volts is impressed across it?


with solution thank you!


1
Expert's answer
2022-04-27T13:34:23-0400

A.

"R=\frac{V}{I}=\rm\frac{120\: V}{12 \: A}=10\: \Omega"

B.

"I=\frac{V}{R}=\rm\frac{50\: V}{100\:\Omega}=0.5\: A"